Let {u,v,w} be an orthornornal set of vectors in R^3. Show that they are linearly independent over R. Check whether u-v, u+v,w are orthogonal over R or not.


1
Expert's answer
2021-02-03T00:11:37-0500

Let "\{u,v,w\}" be an orthornornal set of vectors in "\mathbb R^3". Let us show that they are linearly independent over "\mathbb R". For this consider a linear combination "au+bv+cw=0". Since this vectors are orthornormal, for their scalar products we have that "u\cdot u=v\cdot v=w\cdot w=1,\ \ u\cdot v=v\cdot u=u\cdot w=w\cdot u=v\cdot w=v\cdot w=0".


Therefore, "(au+bv+cw)\cdot u=0\cdot u" implies "a(u\cdot u)+b(v\cdot u)+c(w\cdot u)=0", and thus "a=0." It follows from "(au+bv+cw)\cdot v=0\cdot v" that "a(u\cdot v)+b(v\cdot v)+c(w\cdot v)=0", and thus "b=0." By analogy, "(au+bv+cw)\cdot w=0\cdot w" implies "a(u\cdot w)+b(v\cdot w)+c(w\cdot w)=0", and thus "c=0."

Consequently, the vectors "u,v,w" are linearly independent over "\mathbb R".


Let us show that "u-v, u+v,w" are orthogonal over "\mathbb R". Indeed, their scalar products are equal to zero: "(u-v)\cdot(u+v)=u\cdot u+u\cdot v-v\cdot u-v\cdot v=1+0-0-1=0", "(u-v)\cdot w=u\cdot w - v\cdot w=0-0=0" and "(u+v)\cdot w=u\cdot w + v\cdot w=0+0=0".